# Reduced Rank Regression for Mixed Predictor and Response Variables

### v1

@@ Abstract
In this paper, we propose the generalized mixed reduced rank regression method, GMR3 for short. GMR3 is a regression method for a mix of numeric, binary and ordinal response variables. The predictor variables can be a mix of binary, nominal, ordinal, and numeric variables. For dealing with the categorical predictors we use optimal scaling. A majorization-minimization algorithm is derived for maximum likelihood estimation under a local independence assumption. We discuss in detail model selection for the dimensionality or rank, and the selection of predictor variables. We show an application of GMR3 using the Eurobarometer Surveys data set of 2023.
@@ 1 Introduction
In this manuscript, we will describe a general methodology for the case where we are interested in the analysis of dependence, that is, we like to see the dependence of a set of variables on another set of variables, as in regression methods. Such an analysis of dependence can be contrasted to an analysis of interdependence (Gifi,, 1990), the latter treats the variables in a symmetric way, as in correlation and association analysis.
Usually, researchers have multiple response variables and multiple predictor variables. Typically, researchers analyze the data in a univariate way, one response variable at the time. However, as Fish, (1988) argued, it is important to analyze the data using multivariate methods, as they best honor the reality about which the researcher is purportedly trying to generalize. To quote Fish, (1988), "The reality in which social scientists are interested is usually one in which the researcher cares about multiple outcomes, in which most outcomes have multiple causes, and in which most causes have multiple effects".
Variables can be measured on different scales (Stevens,, 1946). It is possible to distinguish between ratio, interval, ordinal, and nominal variables. Binary variables can be considered special cases of these. In a statistical analysis it is important to take into account the measurement level of the variable, as this defines the kind of operations that are allowed. In the social multivariate context defined above, both the responses as well as the predictors might have varying measurement levels.
@@
These reduced rank models consider one type of response variables, that is, all responses should be numeric, or all should be binary, or all should be ordinal. In empirical research, however, typically the responses have different measurement scales. In the example considered in Section 5, for example, there are binary and ordinal response variables. Luo et al., (2018) developed a reduced rank model for different types of responses. They did consider, numeric, binary and count variables, but not ordinal ones. In the social sciences, ordinal response variables are most common. Furthermore, the different measurement levels on the predictor side are usually taken into account by creating a set of dummy variables. Specifically, for ordinal variables this approach does not take into account the nature of the variable. We will include optimal scaling of the predictor variables.
@@ 2.1 Predictor Variables
The elements of a row of the matrix $\bm{\Phi}$ will be collected in the vector $\bm{\phi}_{i}$, representing the optimally transformed variables for observation $i$ ($i=1,\ldots,N)$.
@@ 2.2 Response Variables
For identification, we require the matrix $\bm{V}$ to be orthogonal, that is $\bm{V}^{\prime}\bm{V}=\bm{I}$, where $\bm{I}$ is the identity matrix of order $S$.
For numeric responses, the $\theta_{ir}$ represents the expected or estimated value of the response $r$. Similarly, for binary response variables the $\theta_{ir}$ represent the log-odds form, that is
@@
For ordinal variables, the story is a bit different. The number of categories of response variable $r$ is $C_{r}$, coded as $c=1,\ldots,C_{r}$. Underlying each ordered categorical response variable $y_{r}$ we assume a continuous latent variable $y^{*}_{r}$. These latent variables are modelled as
@@
We see that for ordinal response variables the $m_{r}=0$, because, without loss of generality, we can assume the latent underlying continuous response variable to be centered, so that no intercept is needed. The continuous underlying variable is partitioned through a set of cut-points or thresholds to form a set of ordered categories. Let $-\infty=t_{0}<t_{1}<\ldots<t_{C_{r}}=\infty$ define the set of thresholds such that an observed ordinal response $y_{r}$ satisfies
@@
where, similar to the proportional odds regression model (McCullagh,, 1980; Anderson and Philips,, 1981), the thresholds are category specific but the structural part of the model is variable specific. From this model specification for ordinal variables, it follows that the probability that person $i$ will respond with category $c$ on response variable $r$ is
@@
a probability that we need in defining the log-likelihood function.
@@ 3 Algorithm
The general algorithm will alternate between updating the canonical part and the threshold parameters. When updating the canonical parameters, we assume the threshold parameters fixed, and vice versa. To avoid cluttering of notation, we will write $\mathcal{L}(\bm{\theta})=\mathcal{L}(\bm{\theta},\bm{t})$ when we update the canonical parameter and $\mathcal{L}(\bm{t})=\mathcal{L}(\bm{\theta},\bm{t})$ when updating the thresholds.
@@ 3.1 MM algorithm for canonical part
where $z=(\vartheta-\frac{1}{\kappa}\xi)$, a working response and $c=c_{1}-\frac{\kappa}{2}z^{2}=\mathcal{L}(\vartheta)-\xi\vartheta+\frac{\kappa}{2}\vartheta^{2}-\frac{\kappa}{2}z^{2}$, all constants with respect to $\theta$. The last line shows that the majorization function is a least squares function. For the different types of response variables, we have to derive the expression of the first derivatives $\xi$ and the majorizing constant $\kappa$.
@@
For ordinal variables, we start out a bit different because we deal with a latent variable and therefore aim for an EM algorithm. In the EM algorithm, the first step is to define the complete data negative log-likelihood, that is, the likelihood assuming we have observed the underlying latent variable. In the E-step, the expected value of this complete data negative log-likelihood is obtained, which in the M-step is minimized. For the minimization, we use an upper bound again like in the MM algorithm.
@@
Let us define $p_{ir}=\frac{1}{1+\exp(-y^{*}_{ir}+\theta_{ir})}$ such that $\log f(y^{*}_{ir}-\theta_{ir})=\log p_{ir}(1-p_{ir})$. The partial derivative is
@@ 3.1.4 Update of Residual Variance
For updating $\sigma^{2}$, we only need the numeric responses. We focus therefore only on the columns related to $r\in\mathcal{N}$. For these columns, we compute the residuals
$\tilde{\bm{Z}}-\bm{1m}^{\prime}-\bm{\Phi}\bm{BV}^{\prime},$
for which we compute the variance to obtain an update of $\sigma^{2}$.
@@ 3.1.5 Update of the Quantifications
For categorical predictor variables, we optimally scale the levels. For numeric variables, the quantified variable is simply the standardized predictor, as discussed in Section 2, that remains constant throughout the iterations. The relevant part of the majorization function for the transformations is
@@
For ordinal predictor variables, in a second step this update is projected onto the cone of admissible quantifications (Meulman et al.,, 2019). This amounts to a weighted monotone regression (De Leeuw,, 2005; Busing,, 2022). Finally, we rescale $\bm{w}_{p}$ such that its mean is equal to zero and its variance equal to one.
@@ 3.2 Estimation of Thresholds
The thresholds of the ordinal response variables are not part of the canonical parameters. Where above we focused on minimizing $\mathcal{L}(\bm{\theta})$, that is, $\mathcal{L}(\bm{\theta},\bm{t})$ with $\bm{t}$ fixed, now we will focus on minimizing $\mathcal{L}(\bm{t})$, that is, $\mathcal{L}(\bm{\theta},\bm{t})$ with $\bm{\theta}$ fixed. Because of the local independence assumption, we can estimate the thresholds separately for each response variable. For estimation of the thresholds for response variable $r\in\mathcal{O}$, we estimate a standard proportional odds model by maximum likelihood where we use the canonical parameter $\theta_{ir}$ as an offset, that is a predictor variable with regression coefficient fixed to 1 (Agresti,, 2013).
4 Model Selection
@@ 5.2.1 Goodness of Fit
Model selection criteria can be based on goodness of fit measures. Both Akaike and Bayesian Information criteria AIC and BIC include the penalization factor accounted for the model complexity as measured by the number of parameters to estimate. For this purpose, Mc Fadden R-squared measure $R^{2}$ has also the adjusted version $R_{a}^{2}$. All these measures are evaluated for all models with an increasing dimensionality $S$. With larger dimensionality, the number ${\cal K}$ of parameters to estimate also increases. Model selection results are shown in Table 1. The optimal choice is for the lowest AIC and BIC as well as the highest $R_{a}^{2}$. In our application, the BIC and $R_{a}^{2}$ agree on the reduced-rank $S=2$ with ${\cal K}=46$ parameters whereas, as often is the case, the AIC proposes a less parsimonious model of reduced-rank $S=3$ with ${\cal K}=53$ parameters.
@@ 5.2.2 Cross validation
From the goodness of fit analysis and the cross validation results, we conclude that the two-dimensional model is the optimal one.

### v2

@@ Abstract
In this paper, we propose the generalized mixed reduced rank regression method, GMR3 for short. GMR3 is a regression method for a mix of numeric, binary, and ordinal response variables. The predictor variables can be a mix of binary, nominal, ordinal, and numeric variables. For dealing with the categorical predictors we use optimal scaling. A majorization-minimization algorithm is derived for maximum likelihood estimation under a local independence assumption. A series of simulation studies is shown (Section 4) to evaluate the performance of the algorithm with different types of predictor and response variables. In Section 5.2, we briefly discuss the choices to make when applying the model the empirical data and give suggestions for supporting such choices. In Section 6.1, we show an application of GMR3 using the Eurobarometer Surveys data set of 2023.
@@ 1 Introduction
In this paper, we will describe a general methodology for the case where we are interested in the analysis of dependence, that is, we like to see the dependence of a set of variables on another set of variables, as in regression methods. Such an analysis of dependence can be contrasted to an analysis of interdependence (Gifi,, 1990), the latter treats the variables in a symmetric way, as in correlation and association analysis.
Usually, researchers have multiple response variables and multiple predictor variables. Typically, researchers analyze the data in a univariate way, one response variable at a time. However, as Fish, (1988) argued, it is important to analyze the data using multivariate methods, as they best honor the reality about which the researcher is purportedly trying to generalize. To quote Fish, (1988), "The reality in which social scientists are interested is usually one in which the researcher cares about multiple outcomes, in which most outcomes have multiple causes, and in which most causes have multiple effects".
Variables can be measured on different scales (Stevens,, 1946). It is possible to distinguish between ratio, interval, ordinal, and nominal variables. Binary variables can be considered special cases of these. In a statistical analysis, it is important to take into account the measurement level of the variable, as this defines the kind of operations that are allowed. In the multivariate social context defined above, both the responses and the predictors may have varying measurement levels.
@@
These reduced rank models consider one type of response variables, that is, all responses should be numeric, or all should be binary, or all should be ordinal. In empirical research, however, typically the responses have different measurement scales. In the example considered in Section 6, for example, there are binary and ordinal response variables. Luo et al., (2018) developed a reduced rank model for different types of responses. They did consider numeric, binary and count variables, but not ordinal ones. In the social sciences, ordinal response variables are most common. Furthermore, the different measurement levels on the predictor side are usually taken into account by creating a set of dummy variables. Specifically, for ordinal variables this approach does not take into account the nature of the variable. We will include optimal scaling of the predictor variables.
@@ 2.1 Predictor Variables
The elements of a row of the matrix $\bm{\Phi}$ will be collected in the $P$-dimensional column vector $\bm{\phi}_{i}$, representing the optimally transformed variables for observation $i$ ($i=1,\ldots,N)$.
@@ 2.2 Response Variables
For identification, we require the matrix $\bm{V}$ to be orthogonal, that is $\bm{V}^{\prime}\bm{V}=\bm{I}$, where $\bm{I}$ is the identity matrix of order $S$. Furthermore, we require $\bm{U}^{\prime}\bm{U}$ to be a diagonal matrix, where $\bm{U}=\bm{\Phi}\bm{B}$.
For numeric responses, $\theta_{ir}$ represents the expected or estimated value of the response $r$. We assume a normal distribution for the numeric response variables.
Similarly, for binary response variables, $\theta_{ir}$ represent the log-odds form, that is,
@@
We assume the binary response variables to have a Bernoulli distribution with probability $pi_{ir}$.
For ordinal variables, the story is a bit different. The number of categories of response variable $r$ is $C_{r}$, coded as $c=1,\ldots,C_{r}$. Underlying each ordered categorical response variable $y_{r}$ we assume a continuous latent variable $y^{*}_{r}$. These latent variables are modeled as
@@
We see that for ordinal response variables the $m_{r}=0$, because, without loss of generality, we can assume that the latent underlying continuous response variable is centered, so no intercept is needed. The continuous underlying variable is partitioned through a set of cut-points or thresholds to form a set of ordered categories. Let $-\infty=t_{0}<t_{1}<\ldots<t_{C_{r}}=\infty$ define the set of thresholds such that an observed ordinal response $y_{r}$ satisfies
@@
where, similar to the proportional odds regression model (McCullagh,, 1980; Anderson and Philips,, 1981), the thresholds are category specific but the structural part of the model is variable specific. From this specification of the model for ordinal variables, it follows that the probability that person $i$ will respond with category $c$ on response variable $r$ is
@@
a probability that we need to define the log-likelihood function. The ordered response follows a multinomial distribution with probabilities $\pi_{irc}$.
@@ 3 Algorithm
The general algorithm will alternate between updating the canonical part and the threshold parameters. When updating the canonical parameters, we assume that the threshold parameters are fixed, and vice versa. To avoid cluttering of notation, we will write $\mathcal{L}(\bm{\theta})=\mathcal{L}(\bm{\theta},\bm{t})$ when we update the canonical parameter and $\mathcal{L}(\bm{t})=\mathcal{L}(\bm{\theta},\bm{t})$ when updating the thresholds.
@@
MM algorithms have several properties. The first is that it usually has a simple numerical minimization method. We will see shortly that in our case the majorization function is a least-squares problem that is relatively easy to solve. Furthermore, the value of the loss function (i.e., in our case the negative log-likelihood) should never increase, which makes it easy to check the programming. MM algorithms are globally convergent and usually end in a local minimum. The disadvantages of MM algorithms are that they are slow (linear convergence rate) and it may be difficult to prove convergence of the parameters.
@@ 3.1 MM algorithm for canonical part
where $z=(\vartheta-\frac{1}{\kappa}\xi)$, a working response and $c=c_{1}-\frac{\kappa}{2}z^{2}$ and $c_{1}=\mathcal{L}(\vartheta)-\xi\vartheta+\frac{\kappa}{2}\vartheta^{2}$, all constants with respect to $\theta$. The last line shows that the majorization function is a least-squares function. For the different types of response variables, we have to derive the expression of the first derivatives $\xi$ and the majorizing constant $\kappa$.
@@
For ordinal variables, we start out a bit different because we deal with a latent variable and therefore aim for an EM algorithm. In the EM algorithm, the first step is to define the complete data negative log-likelihood, that is, the likelihood assuming that we have observed the underlying latent variable. In the E-step, the expected value of this complete data negative log-likelihood is obtained, which in the M-step is minimized. For minimization, we use an upper bound again like in the MM algorithm.
@@
Let us define $p_{ir}=\frac{1}{1+\exp(-y^{*}_{ir}+\theta_{ir})}$ so that $\log f(y^{*}_{ir}-\theta_{ir})=\log p_{ir}(1-p_{ir})$. The partial derivative is
@@
The choice of $\kappa^{*}=\max(\frac{1}{4},\sigma^{-2})$ works fine except when the variance $\sigma^{2}$ is close to zero. As we do not know the value of this variance, we estimate it from the data in each iteration of the algorithm (see Section 3.1.4) and plug the estimate in the algorithm. An estimate close to zero happens, when the numeric response variable are approximated very well. In that case, $\kappa^{*}$ becomes a very large value and the working responses stay very close to the expected value of the previous iteration. In that case, the algorithm gets stuck. In our software, we implemented a warning when the estimated variance becomes very small, that is $\hat{\sigma}^{2}<0.05$.
@@ 3.1.4 Update of Residual Variance
For updating $\sigma^{2}$, we only need the numeric responses. We therefore focus only on the columns related to $r\in\mathcal{N}$. For these columns, we compute the residuals
$\bm{E}=\tilde{\bm{Z}}-\bm{1m}^{\prime}-\bm{\Phi}\bm{BV}^{\prime},$
for which we compute the variance to obtain an update of $\sigma^{2}$ as
$\frac{1}{N\cdot R_{\mathcal{N}}-1}\sum_{i=1}^{N}\sum_{r\in\mathcal{N}}e^{2}_{ir},$
where $R_{\mathcal{N}}$ is the number of numeric response variables.
@@ 3.1.5 Update of the Quantifications
For categorical predictor variables, we optimally scale the levels. For numeric variables, the quantified variable is simply the standardized predictor, as discussed in Section 2, which remains constant throughout iterations. The relevant part of the majorization function for the transformations is
@@
For ordinal predictor variables, we need an extra step. In this extra step this update is projected onto the cone of admissible quantifications (Meulman et al.,, 2019), i.e., the quantifications should have the correct order either increasing or decreasing. This amounts to a weighted monotone regression (De Leeuw,, 2005; Busing,, 2022) of the $\bm{w}_{p}^{+}$ with weights equal to the observed frequencies of each of the response categories. Because the relationship can either be monotone increasing or monotone decreasing, we perform two of such monotone regressions and select the one that fits best. Note that these monotone regression problems are very small, as they only involve the number of categories of the ordered predictor variable.
In the last step of the optimal scaling process, we rescale $\bm{w}_{p}$ such that the mean of $phi_{p}$ is equal to zero and its variance equal to one. This is important, as we multiply the optimally scaled variable with a regression weight and without standardizing we would not be able to obtain unique estimates.
@@ 3.2 Estimation of Thresholds
The thresholds of the ordinal response variables are not part of the canonical parameters. Where above we focused on minimizing $\mathcal{L}(\bm{\theta})$, that is, $\mathcal{L}(\bm{\theta},\bm{t})$ with $\bm{t}$ fixed, now we will focus on minimizing $\mathcal{L}(\bm{t})$, that is, $\mathcal{L}(\bm{\theta},\bm{t})$ with $\bm{\theta}$ fixed. Because of the local independence assumption, we can estimate the thresholds separately for each response variable. For estimation of the thresholds for response variable $r\in\mathcal{O}$, we use standard maximum likelihood estimation where we fixed the other estimates.
3.3 Summary
To summarize our algorithm. The algorithm needs starting values, that we derive using a reduced rank regression model where we take all predictors and responses as numeric variables. Starting with these estimates, we take the following steps:
  • Compute the new working response variables;
  • Update the quantifications ($\phi_{p}$);
  • Update the weights ($\bm{B})$
  • Update the loadings ($\bm{V}$)
  • Update the intercepts for numeric and binary responses ($m$);
  • Update the threshold ($\bm{t}$).
These steps are repeated until the decrease of the negative log-likelihood is smaller than a pre-set convergence criterion (e.g., $1\times 10^{-6}$). The algorithm results in a unique solution. The order of the steps described above can be changed, that would lead to the same estimates.
4 Simulation Study
To evaluate our algorithm, we conducted a simulation study. As discussed in the introduction, Luo et al., (2018) proposed a reduced rank model for mixed numeric, binary, and count variables response variables and numeric predictors. We propose a reduced rank model for mixed numeric, binary and ordinal response variables and mixed predictor variables. We can therefore compare the algorithms for a mixture of binary and numeric response variables with numeric predictors.
We followed the set-up of Luo et al., (2018). We consider their Model 1, a low-dimensional example, with a few adaptions. We set $N=500$, $P=8$, $R=8$, and $S=2$. Among the responses, 4 of them are generated from Normal distribution and 4 from the Bernoulli distribution. The predictor matrix is constructed by generating its entries as independent and identically distributed random samples from the standard normal distribution $\mathcal{N}(0,1)$. The coefficient matrix $\bm{B}$ is an orthogonal matrix from the QR decomposition of a random matrix filled with $\mathcal{N}(0,1)$ entries, and all entries in $\bm{V}$ are samples from the uniform distribution $\mathcal{U}(-1,1)$. We set the intercept vector equal to $\bm{0}$. The numeric responses are drawn from the normal distribution with mean equal to the canonical term and variance equal to one. The binary responses are drawn from the Bernoulli distribution.
To evaluate the performance we compare the population $\bm{BV}^{\prime}$ with the estimated $\hat{\bm{B}}\hat{\bm{V}}^{\prime}$ using the root mean squared error metric (Rmse). The number of replications equals 250. On each generated data set, we fitted the reduced rank model with the algorithm of Luo et al., (2018) and our algorithm. For each algorithm, we thus obtain 250 Rmse measures. We can compare the recovery of both algorithms using boxplots.
Our methodology allows for categorical predictors and ordinal responses. To test our algorithm under these circumstances, we added three conditions. First, considering ordinal predictor variables, we discretized the numeric predictors into five categories based on the .2, .4, .6, and .8 quantiles. For generating the data, we used the average value of the numeric predictors within each of these categories, whereas for data analysis we coded the predictors using the integers 1 to 5.
Second, considering ordinal responses, we generated two different types of data. In the first scenario, 4 binary and 4 ordinal responses were drawn, whereas in the second scenario four numeric and four ordinal variables were drawn. The ordinal response variables have four categories, and were generated by drawing from a multinomial distribution. The estimated thresholds used for deriving the probabilities were set to -1, 0, and 1.
Also for these three extra conditions, we compare the population $\bm{BV}^{\prime}$ with the corresponding estimates using the root mean squared error. We depict the Rmse’s from the 250 replications using a boxplot.
The results are shown in Figure 1. We see that our algorithm performs the same as the algorithm of Luo et al., (2018). Furthermore, we can conclude that changing the predictors into ordinal variables with the use of optimal scaling in the algorithm does not change the performance. The latter two boxplots, show the results with ordinal response variables. When the response variables are a mixture of binary and ordinal variables, the recovery is a bit worse, but still good, whereas the recovery is a little better – compared to a mixture of numeric and binary – when the responses constitute a mixture of numeric and ordinal variables.
Figure 1: Results of simulation study. Different methods are compared in terms of Root mean squared error (Rmse). The first method is the mixed reduced rank regression algorithm of Luo and colleagues, the second is our algorithm. These two are applied to data with numeric predictors and a mix of numeric and binary responses. In the third (gmr3-p), we changed the predictors to be ordinal variables, so that the optimal scaling is employed. In the fourth and fifth method, we generated ordinal responses: in gmr3-r1 the responses are a mixture of binary and ordinal responses, whereas in gmr3-r2 they are a mixture of numeric and ordinal responses.
5 Model Selection for GMR3
Before we show an application on empirical data, we like to set out our thoughts on model selection for our generalized mixed reduced rank regression model. We discuss several existing methods and how they can be used in our framework.
@@ 6.2.1 Goodness of Fit
Model selection criteria can be based on goodness of fit measures. Both Akaike and Bayesian Information criteria AIC and BIC include the penalization factor accounted for the model complexity as measured by the number of parameters to estimate. For this purpose, Mc Fadden R-squared measure $R^{2}$ has also the adjusted version $R_{a}^{2}$. All these measures are evaluated for all models with an increasing dimensionality $S$. With larger dimensionality, the number ${\cal K}$ of parameters to estimate also increases. Model selection results are shown in Table 1. In our application, the BIC suggests a two-dimensional (i.e., $S=2$) model with ${\cal K}=46$ parameters whereas, as often is the case, the AIC proposes a less parsimonious three dimensional model with ${\cal K}=53$ parameters. The adjusted Mc Fadden R-squared measure also suggests three dimensional model, although the differences between the 2, 3, 4, and 5 dimensional models are small.
@@ 6.2.2 Cross validation
From the goodness of fit analysis and the cross validation results, we conclude that the two-dimensional model provides a good representation of the data.
@@ 6.3 Interpretation
6.4 Comparison to standard regression models for each of the responses
We also analyzed the data by running standard regression methods for each of the responses separately. For Trust in the European Union (T) and further expansion of it (FE), this amounts to fitting binary logistic models and for the others (CI. MW, FS, DI, and RE) proportional odds regression models. For the categorical predictor variables we used dummy coding as is usual, with the first category as baseline. We estimate the models and use a bootstrap procedure for finding the standard errors of the parameters. To make a fair comparison, or our GMR3 model, we derived coefficients similar in meaning to those of the separate models. The results are described in detail in Appendix A. We describe here some general conclusions.
Although the total negative log-likelihood is lower than the one obtained with our methodology the sum of AIC and BIC statistics is higher than the AIC and BIC we obtained. This is mainly because of the number of parameters, which for the separate models equals 115, while for our model is 46. The estimated coefficients of the separate models do not take into account the ordered nature of the predictors (see Table 2 and compare to Table 3). Some of the estimated coefficients become really large in the separate models, for example, the effects of the dummy variables for eduction on the response variables MW, FS, DI, and RE. Such large estimates usually point to very unstable results. The implied corresponding estimates of our model are much better. We can also witness the instability of the estimates in the bootstrap standard errors, shown in Table 4. Some of these standard errors are really huge. As a comparison, we showed the standard errors of the corresponding estimates from our model in Table 5. By fitting one low rank model instead of separate regression model, we obtain better estimates that are much stabler. Furthermore, from our results in Figure 3 we directly see that the effects from the predictors on, for example, FS and RE are very similar where such a conclusion is very difficult to obtain from the separate fits.